Find the value of a for which the polynomial x^5-ax²-ax+1has −1 as a root with multiplicity at least 2.
1
Expert's answer
2020-03-04T19:04:34-0500

Solution

If -1 is a root with a multiplicity of at least 2, then this polynomial must be divisible by "(x+1)^2"

"(x+1)^2=x^2+2x+1",

"(x^5-ax-ax+1)/(x^2+2x+1)=\\=x^3-2x^2+3x+\\(-(a+4)x^2-(3+a)+1)/(x^2+2x+1)"

In order for a polynomial to be divisible by "(x^2+2x+1)" , it is necessary that "(-(a+4)x^2-(3+a)+1)" to be divisible by "(x^2+2x+1)" .

Then "-(a+4)=1" and "-(3+a)=2" , hence "a=-5."

Answer

"a=-5."